A ball of mass 200g dropped from a height of 2m, hits the horizontal floor and rebound to a height h below it's initial position. If the coefficient of restitution is 0.8, calculate the value of h

1
Expert's answer
2022-06-06T08:54:04-0400
"e=\frac{v_f}{v_i}=0.8"

"mgh_f=\frac{mv^2_f}{2}=e^2*\frac{mv^2_i}{2}=e^2*mgh_i"

"h_f=e^2h_i=0.8^2*2\: \rm m=1.28\:\rm m"